Data Management Plan for Doron Zeilberger’s NSF proposal Automated Enumerative Combinatorics

1. Expected Data In this project I plan to have numerous Maple packages to perform various sophisticated mathematical tasks. They are all going to be posted in my website, complete with source code, with links from my software page

http://www.math.rutgers.edu/~zeilberg/programs.html .

These programs would produce combinatorial data that might take several days, and sometimes weeks, to produce ab initio. These will be posted as appendices to the front of the relevant articles that would be written.

2. Data Format Source code (written in Maple, and possibly SAGE), as text files. The computer output would be posted as text files.

3. Access to Data and Data Sharing All my data (both software and output) is freely available on my website. There are no restrictions.

4. Policies for Re-use N/A .

5. Archiving of Data Everything is on my website at Rutgers University, with daily backups. In addition I have backup copies of all my software at my home computer.